Exercise 2 (i If is open in C and Ac C any subset, then ON A is open in A (ii) If A C C is any subset and U is open in A, thethis question is about complex variables
Answer #1

(i). Given \Omega is a open set in \mathbb{C} and AC be any set .

Let E A be an arbitrary element ,

E2, E A

As ES and \Omega is open so there exist a neighbourhood of x say Nx such that x \in N_{x} \subset \Omega

NnAC A is an neighbourhood of x containted in \Omega \cap A .

\Rightarrow x is an interior point of \Omega \cap A , sincce x is arbitrary so every point of \Omega \cap A is an interior point .

Hence \Omega \cap A is open in A .

(ii). A \subset \mathbb{C} be any set and U is open in A .

Let , \Omega = int (A) \cap U i.e., \Omega is the collection of all interior point of A which is an U so \Omega is an open set as it is intersection of two open set .

As U a open subset of A so each point of U is an interior point of A

\RightarrowU \subset int(A)

Hence ,

\Omega \cap A

= int (A) \cap U\cap A

= U , since U \subset A , U\subset int(A) .